OpenStudy (anonymous):

Please help! @Astrophysics @DanJS Luke buys a certain brand of cereal that costs $11 per box. Luke changes to a super-saving brand of the same size. The equation shows the price, y, as a function of the number of boxes, x, for the new brand. y = 9x Part A: How many more dollars is the price of a box of Luke’s original brand of cereal than the price of a box of the super-saving brand? Show your work. Part B: How much money does Luke save each month with the change in cereal brand if he buys 6 cereal boxes each month? Show your work.

OpenStudy (ali2x2):

For part A, we know that the new box is $9 per box, and his original was $11 per box, so subtract 11 - 9 to get the answer. So what is your answer for part A?

OpenStudy (ali2x2):

@kool_22

OpenStudy (anonymous):

11-9=2

OpenStudy (ali2x2):

good job :)

OpenStudy (ali2x2):

If he buys 6 boxes each month, you multiply 11 x 6 for the original box, and 9 x 6 for the new one.

OpenStudy (ali2x2):

@kool_22

OpenStudy (ali2x2):

if you dont know just ask and ill tell you :)

OpenStudy (anonymous):

11*6= 66 9*6= 54 and sorry I keep leaving my computer only lets me type in 1 answer at a time. and it is very slow when it comes to openstudy

OpenStudy (ali2x2):

Huh ok well thats fine :) And yes your correct

OpenStudy (ali2x2):

Subtract 66 - 54

OpenStudy (anonymous):

66-54= 12

OpenStudy (ali2x2):

Thus having your answer :)

OpenStudy (ali2x2):

Yes, so he will save $12 each month if he buys 6 boxes a month.